# Breaking the News: A LLM-based Game where Players Act as Influencer or Debunker for Raising Awareness About Misinformation

### v1

@@ Abstract.
Game-based interventions are widely used to combat misinformation online by employing the ”inoculation approach”. However, most current interventions are designed as single-player games, presenting players with limited predefined choices. Such restrictions reduce replayability and may lead to an overly simplistic understanding of the processes of misinformation phenomenon and the debunking. This study seeks to address these issues, and empower people to better understand the opinion influencing and misinformation debunking processes. We did this by creating a Player versus Player (PvP) game where participants attempt to either generate or debunk misinformation to convince LLM-represented public opinion. Using a within-subjects mixed-methods study design (N=47), we found that this game significantly raised participants’ media literacy and improved their ability to identify misinformation. Our qualitative exploration revealed how participants’ use of debunking and content creation strategies deepened their understanding of the nature of disinformation. We demonstrate how LLMs can be integrated into PvP games to foster greater understanding of contrasting viewpoints and highlight social challenges.
1, 2, 3
journalyear: 2023copyright: acmlicensedconference: Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems; April 23–28, 2023; Hamburg, Germanybooktitle: Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23), April 23–28, 2023, Hamburg, Germanyprice: 15.00doi: 10.1145/3544548.3581464isbn: 978-1-4503-9421-5/23/04ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
The prevalence of misinformation is a growing concern globally. Misinformation damages society in numerous ways. It threatens the maintenance of trust in vaccines and health policies(Do Nascimento et al., 2022; MacDonald and Brianna, 2023), can incite violence and harassment(Sinders, 2022), can undermine democratic processes (particularly elections)(Bovet and Makse, 2019; Groshek and Koc-Michalska, 2017), and harms individual and societal well-being(Verma et al., 2022). For example, during the 2018 Brazilian presidential election, manipulated photos, decontextualized videos, and hoax audio significantly influenced election results by aiding the victory of the main far-right candidate (Avelar, 2019; Santos, 2020). Another example is the spread of false information during the COVID-19 pandemic. Many people were persuaded that ineffective or counterproductive treatments such as alcohol-based cleaning products and an anti-parasitic drug could cure patients. The victims of this misinformation could suffer serious illness or even death (Coleman, 2020).
Different countermeasures have been investigated to combat misinformation. These interventions can be categorized into two major groups: pre-emptive intervention (”prebunking”) and reactive intervention (”debunking”)(Ecker et al., 2022). Debunking involves fact-checking and correcting misinformation after it has been encountered. Fact-checkers use investigative practices to determine the veracity of content and dispute factual inaccuracies(Juneja and Mitra, 2022).
However, the lasting effects of misinformation make it challenging to mitigate its influence once people have been exposed (Roets et al., 2017; Lewandowsky et al., 2012). Furthermore, fact-checking efforts are limited in terms of scale and reach, which restricts their effectiveness(Roozenbeek et al., 2020b). Given these challenges, it is not sufficient to rely purely on debunking efforts. Prebunking, on the other hand, works to build attitudinal inoculation, enabling people to identify and resist manipulative messages. This approach equips individuals to better manage misinformation they encounter in the real world(Soni, 2024). Prebunking is based on a range of educational measures (Dame Adjin-Tettey, 2022; Cook et al., 2017), which can include games (Roozenbeek and Van der Linden, 2019; Cook et al., 2023).
Recent research indicates that the game-based learning approaches are potentially useful prebunking interventions. These games educate and build resistance in people exposed to misinformation(Traberg et al., 2022; Kiili et al., 2024). In games such as ”Bad News,”(Roozenbeek and Van der Linden, 2019) ”Harmony Square,”(har, 2020) ”Go Viral!”(Lewsey, 2020) and ”Trustme!”(Yang et al., 2021) players adopt the role of a misinformation producer whose task is to create and spread misinformation as efficiently as possible. Another approach, applied in the games ”MAthE”(Katsaounidou et al., 2019) and ”Escape Fake”(esc, 2022), involves players acting as fact-checkers and challenging them to assess the validity of information. In some scenarios players can use tools such as search engines to gather clues.
While gamified misinformation interventions have shown promise in teaching players about the nature of misinformation, the choice-based formats of these games can limit replayability. Additionally, such choice-based game formats require little cognitive effort from the player, who is presented with a limited number of pre-generated options. This can diminish player involvement in the game and reduce enjoyment of gameplay. Another factor limiting engagement is that most of these games are designed for single-player mode. Multiplayer mechanics, by contrast, are an effective way to enhance motivation and replayability in games(Mustaro and Mendonça, 2012). Social interactions also significantly contribute to player engagement in educational and other serious games(Lepper and Malone, 2021). Research has demonstrated that both collaborative and competitive gameplay can enhance the effectiveness of serious games(Cagiltay et al., 2015; Bellotti et al., 2010).
@@
To address the challenges of non-natural interactions and a deterministic game paths (with a view to fostering more open-ended exploration and engagement through Player versus Player (PvP)), we need to incorporate more complex scenarios. These would give players the ability to not only choose from preselected options, but also to actively generate content and implement their own strategies for debunking or creating misinformation. Additionally, implementing a PvP approach can be a training tool through which players can learn how to counter real-life misinformation.
@@
Progress in AI and Natural Language Processing (NLP) provide new opportunities for creating more engaging game experiences. Large language models (LLMs) can simulate complex human interactions and societal dynamics(Ziems et al., 2024)and have been applied in video games for tasks like generating narratives(Park et al., 2023a), non-playable characters (NPCs),dialogue(Ashby et al., 2023; Uludağlı and Oğuz, 2023), and role-playing scenarios(Xu et al., 2023). Previous work also demonstrated the opportunities for prompting the LLM to impersonate a specific character and create interactive dialogues from this perspective(Zhou et al., 2024; Shao et al., 2023). Integrating LLMs into the game can introduce greater variability to in-game interactions and enhances both engagement and replayability. As the effectiveness of many inoculation interventions tends to diminish over time, developing an enjoyable and replayable game that consistently reinforces players’ cognitive ”resilience” against misinformation is an essential advancement in this field(Wells, 2024). Applying LLM to the game also has the potential to increase the educational impact of the intervention. It allows users not just to select from a range of options but also to put their input into the model and receive individual feedback tailored to this input.
@@
Inspired by previous misinformation game interventions, this research has involved the development of a PvP misinformation education game called Breaking the News. In our game, two players are assigned either the role of a misinformation creator (referred to as an ”influencer” in the game) or a counteractor of misinformation (referred to as a ”journalist-debunker” in the game). The influencer creates misinformation posts in a system that mimics a social media environment, while the debunker seeks to counter these messages by presenting compelling arguments. LLMs are used to represent public opinion in the ”country” where the game events take place. The goal for both players is to earn the trust of the citizens and convince them to believe the information they present.
In this paper we aim to answer the following research questions:
@@
RQ1: How may we empower players to understand the processes of misinformation generation and misinformation debunking through a GenAI-based PvP game?
RQ2: What behaviors do players exhibit when they are asked to generate versus protect against misinformation?
@@
In this paper, we present the design and evaluation of the PvP game. We conducted a mixed-method study with 47 participants, using a within-subjects design and pre- and post-surveys for repeated measures. Our findings suggest that through gameplay, participants improved their ability to reflect on instances of misinformation, raised their levels of media literacy, expanded their repertoire of strategies applied to countering misinformation, and improved their discernment abilities. This study contributes to the growing body of work analyzing misinformation education games. Specifically, we provide insights into integrating LLMs and interactive PvP mechanics in media literacy contexts. We also offer practical guidance for the design of serious games aimed at combating misinformation in a dynamic, real-world manner.
@@ 2.1. Characteristics of Misinformation
The term ”misinformation” is often used to refer to false information, including so-called ”fake news”, falsehoods, malicious rumours, and conspiracy theories. Some scholars distinguish between misinformation and disinformation, with the latter referring to information deliberately crafted and spread with the intent to deceive or cause harm(Guess and Lyons, 2020; Swire-Thompson et al., 2020). However, as we believe it is often difficult to prove intent (a view inspired by recent publications in the field) this paper uses ”misinformation” as an umbrella term for all forms of false information(Ecker et al., 2022; Southwell et al., 2019; Wu et al., 2019; MIT Generative AI, 2023; Hartwig et al., 2024).
@@
Several studies have sought to identify the key characteristics that distinguish misinformation from well-sourced, authentic information. This analysis aims to helppeople to better make this distinction. For example, Horne et al. analyzed the writing styles of fake information versus real news. They found that texts which can be characterized as fake news typically features longer headlines, simpler word choices, and greater use of proper nouns and verb phrases(Horne and Adali, 2017). Research has shown that misinformation and authentic information differ in terms of how the former can be created with the intention of triggering emotions such as fear, anxiety, or sympathy(Choudhary and Arora, 2021; Zhang and Ghorbani, 2020). In order to produce these responses, misinformation creators employ personal, negative, and opinionated styles (Bessi et al., 2015; Porat et al., 2019). Source credibility is another notable characteristic of distinguishing between authentic information and misinformation. Authentic information is usually shared by highly credible sources: reputable websites, mainstream media outlets, professional news organizations, and official state publications (Molina et al., 2021; Zhang and Ghorbani, 2020; Shin et al., 2018). In contrast, fake news often originates from lower-quality sources designed to generate revenue through advertisements. To attract clicks, these stories frequently use unverified quotes, inflammatory narratives, and misleading images. (Molina et al., 2021; Shu et al., 2017). However, not all less reliable sources are perceived as equally untrustworthy. People often gravitate toward partisan sources that align with their political ideology, leading to significant variations in news consumption across the political spectrum (Faris et al., 2017). This selective consumption reinforces trust in these sources, even when they may be objectively classified as biased or unreliable (Mummolo, 2016; Pennycook and Rand, 2019). Finally, compared to authentic news, misinformation tends to develop in line with a broad dynamic pattern. Shin et al. traced the life-cycle of 17 high-profile political rumors on Twitter over a 13 month period and found that false rumors tend to reemerge, become more intense and extreme over time(Shin et al., 2018). In combination with the ”illusion of truth effect” (i.e., the more often people see information, the more likely they are to trust it), it makes people more likely to believe increasingly extreme misinformation, thus making it harder to debunk(Silva et al., 2017; Yousif et al., 2019).
2.2. Media Literacy
Media literacy is commonly defined as the ability to ”access, analyze, and produce information for specific outcomes” ((Aufderheide, 2018), p. 6). Most modern frameworks on media literacy interpret the term ”information” as being produced by media (including social media), while noting that media outlets can generate adversarial tensions in consumers. There is understanding of a need in democracies for citizens to learn a set of skills that can help protect societies against these effects (Potter, 2010). There are several models outlining the skills that are required to be learned to achieve media literacy. For example, Tandoc et al. proposed a four-component model, which includes technical competency (instrumental knowledge about how to use social media platforms), privacy protection (knowledge about how to protect privacy and manage account security), social literacy (knowledge about how to manage social relations on various platforms), and information awareness (the ability to discern between truthful and false information on social media, identify reputable sources, and cross-check information)(Tandoc Jr et al., 2021). Chen et al. proposed dividing the skills into critical and functional domains, as well as consuming and producing (presuming) content. In this model, the functional skills are related to the basic competencies required to use and produce media content (e.g., accessing a platform and using integrated tools to write a post). This contrasts with critical skills which involve reflecting on the content, recognizing the motives behind publications, and creating content that includes the author’s perspective.(Chen et al., 2011). There are also related concepts of ”news literacy” (Malik et al., 2013) and ”information literacy”(Jones-Jang et al., 2021), which to some extent can be considered part of the media literacy domain specifically aimed at assessing the quality of information and detecting fake news. As mentioned by Hobbs and Jensen (Hobbs and Jensen, 2009), at the start of the 21st century, Media Literacy Education experienced a paradigm shift from a protectionist position (shielding vulnerable individuals from the negative influence of media) to an empowering paradigm, where people were encouraged to critically engage with media and develop skills to interpret both its negative and positive effects through critical thinking. In this paradigm, interaction with misinformation can also be considered as having educational power if it teaches the person to understand its effects. This has resulted in several recent game-based interventions that use inoculation theory to create media literacy education strategies. Inoculation theory suggests that exposure to a weaker version of misinformation can help to develop stronger protection against future exposure (Grace and Liang, 2023).
@@ 2.3. Serious Games in Media Literacy Domain
In recent years, games have gained significant attention as educational tools, driven by their potential to enhance learning experiences (De Freitas, 2018). Game-based learning and gamification, which integrate gaming elements into education, have been studied for their ability to engage learners and facilitate active, experiential learning. These approaches provide immediate feedback and contextual problem-solving and are supported by theories of effective learning (Plass et al., 2015; Tobias et al., 2014). The concept of serious games was proposed by Michael et. (Michael and Chen, 2005); the most common definition of serious games or applied games is that games that do not primarily aim at entertainment, enjoyment, or fun but explicitly emphasize the added educational value of engagement and competition. Troyer et al. (De Troyer et al., 2017) suggest serious games combine serious educational content with the fascinating mechanics, storytelling and narrative of video games, thus making them an effective tool for conveying educational messages in an interactive and engaging way. Serious games have a wide range of applications across various educational settings (Connolly et al., 2012, 2008). There are multiple attempts to gamify misinformation education aiming to improve their media and information literacy (Roozenbeek and Van der Linden, 2019). These games feature players that can be categorized into two main types: creators and debunkers. The misinformation creator’s objective is to create and spread misinformation. For example, in Bad News (Roozenbeek and Van der Linden, 2019), Harmony Square (har, 2020), Cat Park (Gusmanson.nl, 2022) and ChamberBreaker (Jeon et al., 2021), players are tasked with spreading fake news in a social media environment to gain likes or followers while maintaining credibility. The second one focuses on fact-checking, where players adopt the role of a misinformation debunker. In games like MathE (Katsaounidou et al., 2019) and Escape the Fake (esc, 2022), players work to identify fake news using verification tools such as reverse image search. These games support the development of practical skills in identifying misinformation and offer a more direct method of engaging with fake news detection.
@@
These games are typically designed as single-player games, allowing players to experience the game narrative at their own pace. One of the most popular examples is Bad News (Roozenbeek and Van der Linden, 2019)- In Bad News, players actively learn the strategies used to create and spread fake news within the game’s narrative, such as the use of emotionally charged content and the manipulation of social media platforms. Roozenbeek and van der Linden (Roozenbeek and Van der Linden, 2019) conducted a large-scale evaluation of this game, providing evidence that through this serious game mechanic, players become more aware of the psychological techniques behind misinformation, thus improving their ability to critically evaluate real-world information. Another notable example is Fakey (Micallef et al., 2021) in which the goal of the player is to support a healthy social media experience by promoting information from reliable sources rather than from low-credibility sources. While the multiplayer approach is less popular, there are several attempts to create PvP/team settings for misinformation games. For example, FakeYou! (Clever et al., 2020) offers a simple mobile environment where players can create fake news headlines and test their ability to spread misinformation by challenging another player’s ability to recognize misinformation. DoomScroll (Wells et al., 2024) proposes a team-based playing mode, where several players together try to tackle misinformation-related challenges.
2.4. AI-Driven Solutions in Combat Misinformation and Educational Games
AI-based technologies are widely used in combating misinformation. These include automated fact-checking (Choi and Ferrara, 2024), AI-based credibility indicators (Lu et al., 2022), AI and LLM-based explanations of content veracity (Horne et al., 2019; MIT Generative AI, 2023), and personalized AI fact-checking systems (Jahanbakhsh et al., 2023). Most of these efforts focus on a debunking approach, where false information is identified and corrected after dissemination. In the context of prebunking and media education, there was an attempt to develop a role-playing game where players acted as fact-checkers, reviewing storylines generated by LLMs. The results showed that this approach could empower players to critically engage with misinformation through investigative role-play (Tang and Singha, 2024). Gallotta et al. analyzed the LLM roles used in game and identified LLM-based agent under specific roles are well-suited for real-time, context-based interaction(Gallotta et al., 2024), which can provide players personalized game experiences. Generative LLM-based agents can simulate human behavior based on past events and reflection (Park et al., 2023b). Wang et al. introduced a system where LLM-based narrative planning and character simulation create dynamic stories(Wang et al., 2024). Similarly, Peng et al. explored a text-adventure game where, under a fixed narrative premise, players can freely interact with NPCs generated by GPT-4, leading to emergent gameplay behaviors(Peng et al., 2024).
Challenges such as so called ”hallucinations” remain when LLM produce false information, as shown in research by Park et al. (Park et al., 2023b). Previous studies also noted memory limitations in these models. While solutions have been discussed, long interactions and excessive prompts come with the risks of inconsistent outputs (Zhong et al., 2024). Also, variations in writing style and spelling in the input text can significantly impact the outputs, sometimes to the extent of producing incoherent outputs (Jessika Rumbelow, 2024).

### v2

# Breaking the News: Taking the Roles of Influencer vs. Journalist in a LLM-Based Game for Raising Misinformation Awareness

@@ Abstract.
Effectively mitigating online misinformation requires understanding of their mechanisms and learning of practical skills for identification and counteraction. Serious games may serve as tools for combating misinformation, teaching players to recognize common misinformation tactics, and improving their skills of discernment. However, current interventions are designed as single-player, choice-based games, which present players with limited predefined choices. Such restrictions reduce replayability and may lead to an overly simplistic understanding of misinformation and how to debunk them. This study seeks to empower people to understand opinion-influencing and misinformation-debunking processes. We created a Player vs. Player (PvP) game in which participants attempt to generate or debunk misinformation to convince the public opinion represented by LLM. Using a within-subjects mixed-methods study design (N=47), we found that this game significantly raised participants’ media literacy and improved their ability to identify misinformation. Qualitative analyses revealed how participants’ use of debunking and content creation strategies deepened their understanding of misinformation. This work shows the potential for illuminating contrasting viewpoints of social issues by LLM-based mechanics in PvP games.
Inoculation, Misinformation, Generative AI, Game-based learning
copyright: ccdoi: 10.1145/3748600journalyear: 2025journal: PACMHCIjournalvolume: 9journalnumber: 6article: GAMES005publicationmonth: 10ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
The prevalence of misinformation on social media is a growing global concern. Misinformation threatens the maintenance of trust in social agendas like vaccine and health policies (Do Nascimento et al., 2022), incites violence and harassment (Sinders, 2022), undermines democratic processes (Bovet and Makse, 2019), and harms individual and societal well-being (Verma et al., 2022). For example, people were even persuaded to take ineffective treatments like alcohol-based cleaning products and anti-parasitic drugs for Covid-19. Countermeasures against misinformation spread consist of two forms: preemptive intervention (“prebunking”) and reactive intervention (“debunking”)(Ecker et al., 2022). The latter involves correcting misinformation after it has been encountered. and using fact-checking to dispute factual inaccuracies. However, the lasting effects of misinformation make it challenging to mitigate its influence once people have been exposed (Roets et al., 2017), and fact-checking efforts are limited in scale and reach (Roozenbeek et al., 2020b). Prebunking, on the other hand, works to build attitudinal inoculation, enabling people to identify and resist manipulative messaging. This approach equips individuals to manage misinformation they encounter in the real world (Soni, 2024) based on educational measures, including games (Cook et al., 2023; Fu et al., 2025).
Game-based prebunking interventions engage users in simulated misinformation scenarios, allowing them to actively practice identifying and countering deceptive content(Kiili et al., 2024). On one hand, in games like “Bad News,”(Roozenbeek and Van der Linden, 2019) “Harmony Square,”(Roozenbeek and van der Linden, 2020) “Go Viral!”(Lewsey, 2020) and “Trustme!”(Yang et al., 2021), players adopt the role of a misinformation producer whose task is to create and spread misinformation as efficiently as possible. In other games, like “MAthE”(Katsaounidou et al., 2019) and “Escape Fake”(Paraschivoiu et al., 2021), players acting as fact-checkers assessing the validity of information. The choice-based formats of these games can limit replayability, requiring little cognitive effort from the player, who is presented with limited numbers of pre-generated options that reduce the involvement of players. These games are also designed for single players without utilizing multiplayer mechanics that can enhance motivation through social play(Lepper and Malone, 2021). Instead, collaborative and competitive gameplay can enhance the effectiveness of serious game interventions(Cagiltay et al., 2015).
@@
To address the challenges of limited interactions and deterministic game paths, we aim to foster open-ended exploration and engagement through Player versus Player (PvP) mechanics. This provides players the ability to not only choose from preselected options, but also to actively create content and implement their own strategies for disproving or creating misinformation as a training ground for countering real-life misinformation.
@@
Recent work in Large language models (LLMs) provide new interaction possibilities for in-game natural dialogue (Zhou et al., 2024b), generating narratives(Park et al., 2023), non-playable characters (NPCs) interactions (Ashby et al., 2023; Uludağlı and Oğuz, 2023; Gao et al., 2024), and role-playing scenarios(Xu et al., 2023). LLMs can be prompted to impersonate specific characters to provide appropriate dialogue (Zhou et al., 2024a; Zhou and Lc, 2025; Zhang et al., 2025). Integrating LLMs into creative processes in the interaction can increase involvement and critical engagement(Yang et al., 2022; Han et al., 2024; Zeng et al., 2025). As the effectiveness of many inoculation interventions tends to diminish over time, developing a replayable game that reinforces players’ cognitive “resilience” against misinformation is an essential(Wells, 2024). Instead of selecting from predefined inputs, applying LLM to interactions enables users to put their input into the model and receive individual feedback tailored to this input.
@@
Inspired by previous misinformation game interventions, we developed a PvP game called Breaking the News, where players are assigned either the role of a misinformation creator (“Influencer” in the game) or a counteractor against misinformation (“Journalist” in the game). The Influencer creates misinformation posts in a social media-like environment, while the Journalist seeks to counter these messages by presenting compelling arguments. LLMs are used to represent public opinion in the “country” where the game events take place. The goal for both players is to earn the trust of the citizens and convince them to believe the information they present. In this paper we aim to answer the following research questions:
@@
RQ1: How may we empower users to recognize and understand the processes of misinformation generation and misinformation debunking?
RQ2: What behaviors do players exhibit in response to game mechanics and opponent tactics?
@@
In this paper, we present the design and evaluation of the PvP game. We conducted a mixed-method study with 47 participants, using within-subjects design and pre- and post-surveys for repeated measures. Our findings suggest that through gameplay, participants improved their ability to reflect on instances of misinformation, raised their levels of media literacy, expanded their repertoire of strategies applied to countering misinformation, and improved their discernment abilities. This study provides insight into applications of LLMs in interactive PvP mechanics for media literacy. We offer practical insights for the design of serious games aimed at combating misinformation in real-world contexts.
@@ 2.1. Characteristics of Misinformation
The term “misinformation” is often used to including “fake news”, falsehoods, malicious rumours, and conspiracy theories. Some scholars distinguish between misinformation and disinformation, with the latter referring to information deliberately crafted and spread with the intent to deceive or cause harm(Guess and Lyons, 2020). Since it is often difficult to prove intent, we use “misinformation” as an umbrella term for diverse forms of false information(Southwell et al., 2019).
@@
Studies have sought to identify the key characteristics that distinguish misinformation from well-sourced, authentic information. One work analyzed the writing styles of fake information versus real news. They found that texts which can be characterized as fake news typically feature longer headlines, simpler word choices, and greater use of proper nouns and verb phrases(Horne and Adali, 2017). Misinformation and authentic information appear to differ in how the former can be created with the intention of triggering emotions such as fear, anxiety, or sympathy(Choudhary and Arora, 2021) using opinionated wording (Porat et al., 2019). Source credibility also differ between authentic information and misinformation. Authentic information is shared by credible sources: reputable websites, mainstream media outlets, professional news organizations, and official publications (Zhang and Ghorbani, 2020). In contrast, fake news often originates from sources designed to generate revenue. To attract clicks, these stories use unverified quotes, inflammatory narratives, and misleading images (Molina et al., 2021).
However, not all less reliable sources are perceived as equally untrustworthy. People often gravitate toward partisan sources that align with their political ideology, leading to variations in news consumption across the political spectrum (Faris et al., 2017). This selective consumption reinforces trust in these sources, even when they may be classified as biased or unreliable (Pennycook and Rand, 2019). Compared to authentic news, misinformation tends to develop in line with a broad dynamic pattern. One study traced the life-cycle of high-profile political rumors on Twitter over a 13 month period and found that false rumors tend to reemerge, becoming more extreme over time(Shin et al., 2018). Since people are more likely to trust information that they see more often, consumers are led to believe increasingly extreme misinformation that are hard to debunk(Yousif et al., 2019).
2.2. Media Literacy as Protection Against Misinformation
Media literacy is defined as the ability to “access, analyze, and produce information for specific outcomes” (Aufderheide, 2018). Citizens often need to learn skills that can protect societies against misinformation (Potter, 2010). One work proposed a four-component model of media literacy, including technical competency, privacy protection, social literacy, and information awareness, the latter defined as the ability to discern between truthful and false information on social media(Tandoc Jr et al., 2021). Chen et al. proposed dividing the skills into critical and functional domains, as well as consuming and producing content. This critical skills often need to be developed for reflecting on the content, recognizing the motives behind publications, and creating content that includes the author’s perspective as forms of social influence (Chen et al., 2011; LC and Mizuno, 2021; LC et al., 2022; Song et al., 2021, 2022).
Media literacy education has recently shifted from a protectionist position to an empowering paradigm, where people were encouraged to critically engage with media and develop skills to interpret its effects (Hobbs and Jensen, 2009). In this paradigm, interaction with misinformation can also be considered as having educational power if it teaches the person to understand its effects. This has resulted in recent game-based interventions that use inoculation theory to enact media literacy education. Inoculation theory suggests that exposure to a weaker version of misinformation can help to develop stronger protection against future exposure (Grace and Liang, 2023).
@@ 2.3. Serious Games in Media Literacy Domain
Game-based learning and gamification, which integrate gaming elements into education, have been studied for their ability to engage learners and facilitate active, experiential learning. These approaches provide interactive feedback and contextual problem-solving, supported by theories of effective learning (Tobias et al., 2014). Serious games or applied games is defined as games primarily aim at educational value of engagement and competition. Serious games may combine educational content with playful mechanics and game narrative for conveying educational messages interactively(De Troyer et al., 2017; LC and Ruijters, 2022; LC et al., 2020), often using speculative elements to push players into game worlds with metaphorical practical social good applications(Gong et al., 2025; Qian et al., 2025).
Serious games have begun to address misinformation education by aiming to improve media literacy (Roozenbeek and Van der Linden, 2019). These games pit players into either creators and fact-checkers. The misinformation creator’s objective is to create and spread misinformation. For example, in Bad News (Roozenbeek and Van der Linden, 2019), Harmony Square (Roozenbeek and van der Linden, 2020), Cat Park (Gusmanson.nl, 2022) and ChamberBreaker (Jeon et al., 2021), players are tasked with spreading fake news in a social media environment to gain likes or followers while maintaining credibility. In the fact-checking role, games like MathE (Katsaounidou et al., 2019) and Escape the Fake (Paraschivoiu et al., 2021) have players work to identify fake news using verification tools such as reverse image search.
@@
One example of such single-player games is Bad News (Roozenbeek and Van der Linden, 2019). Here, players actively learn the strategies used to create and spread fake news within the game’s narrative, such as the use of emotionally charged content and the manipulation of social media platforms. Through these mechanisms, players became aware of the psychological techniques behind misinformation, improving their ability to critically evaluate real-world information. By contrast, in the game Fakey, the goal of the player is to support a healthy social media experience by promoting information from reliable rather than low-credibility sources (Micallef et al., 2021). Similarly, in the game Cranky Uncle, players were shown popular misconceptions related to the vaccine and shown how to counter them(Cook et al., 2023).
There are also work adapting PvP and team mechanics in serious games about misinformation. FakeYou! is a mobile game where players create fake news headlines and test their ability to spread misinformation by challenging another player’s ability to recognize misinformation (Clever et al., 2020). DoomScroll proposes a team mode where players tackle misinformation challenges together (Wells et al., 2024).
2.4. AI-Driven Interactions for Misinformation Education
AI-based technologies currently employed in combating misinformation include automated fact-checking (Choi and Ferrara, 2024), AI-based credibility indicators (Lu et al., 2022), AI and LLM-based explanations of content veracity (Horne et al., 2019), and personalized AI fact-checking systems (Jahanbakhsh et al., 2023). These efforts focus on debunking interventions, where false information is identified and corrected after dissemination. There are fewer works in prebunking forms of media education, although one study empowers players to critically engage with misinformation through investigative role-play (Tang and Singha, 2024).
Recent advances in Generative AI (GenAI) LLM-based agents can simulate human behavior based on past events and reflection (Zhou et al., 2025; He et al., 2025; Liu et al., 2025). These agents can be designed for gaming contexts using natural text-based descriptions (Ling et al., 2024; Ma et al., 2025) to support narrative design (Fu et al., 2024). Applications of these types of GenAI enabled designs include a text-adventure game where players can freely interact with NPCs generated by GPT-4, leading to emergent gameplay behaviors(Peng et al., 2024), and a GPT dialogue-based game in a speculative post-climate world(Zhou et al., 2024b).
However, working with LLMs can lead to “hallucinations” when LLM spontaneously produce false information (Park et al., 2023), as well as bias and stereotype. Attempts to address this include excessively long prompts that come with its own risks of inconsistent outputs (Zhong et al., 2024). Even variations in writing style and spelling in the input text can impact the outputs, leading to generation of incoherent outputs (Han et al., 2024; Chen et al., 2025). These challenges of working with LLMs must be addressed during careful prompting in the game design process.
2.5. Mapping Game Design to Media Literacy Constructs
Prior misinformation game interventions typically adopt single-player, choice-based formats that aim to affect player recognition abilities rather than production competencies. To address this, we aim to design a game to foster both critical consuming and critical prosuming skills through competitive play. We employed a PvP format with free-form responses, enabling players to craft persuasive or corrective arguments based on both game background news and their opponent’s messages. We integrated LLM-simulated evaluators that dynamically assess player messages, thereby improving critical consuming skills through active interpretation of personalized feedback. Together, these design choices aim to support participants’ development of media literacy skills while offering an immersive and replayable learning experience.